# arXiv:2512.09295v2[math.ST] 01 Mar 2026

### v2

@@ Appendix A Technical Proofs
Proof of Lemma 9.
This is a standard result in empirical process theory; see [GZ90], and [Lia21, Lemma 26]. We sketch the proof for completeness. The proof follows by realizing the fact that $\mathop{\mathbb{E}}[g(Y)]=\mathop{\mathbb{E}}_{\{Y_{i}^{\prime}\}_{i=1}^{n}}\widehat{\mathop{\mathbb{E}}}_{n}^{\prime}[g(Y)]$, where $\{Y_{i}^{\prime}\}_{i=1}^{n}$ is an independent copy of $\{Y_{i}\}_{i=1}^{n}$, and thus
$\mathop{\mathbb{E}}_{\{Y_{i}\}_{i=1}^{n}}\sup_{g\in\mathcal{G}}(\mathop{\mathbb{E}}-\widehat{\mathop{\mathbb{E}}}_{n})[g(Y)]\leq\mathop{\mathbb{E}}_{\{Y_{i},Y_{i}^{\prime}\}_{i=1}^{n}}\sup_{g\in\mathcal{G}}(\widehat{\mathop{\mathbb{E}}}_{n}^{\prime}-\widehat{\mathop{\mathbb{E}}}_{n})[g(Y)]$ $=\mathop{\mathbb{E}}_{\{\epsilon_{i}\}_{i=1}^{n}}\mathop{\mathbb{E}}_{\{Y_{i},Y_{i}^{\prime}\}_{i=1}^{n}}\sup_{g\in\mathcal{G}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}(g(Y_{i}^{\prime})-g(Y_{i}))$

### v3

@@ Appendix A Technical Proofs
Proof of Lemma A.9.
This is a standard result in empirical process theory; see [GZ90], and [Lia21, Lemma 26]. We include the argument for completeness. The proof follows from the observation that $\mathop{\mathbb{E}}[g(Y)]=\mathop{\mathbb{E}}_{\{Y_{i}^{\prime}\}_{i=1}^{n}}\widehat{\mathop{\mathbb{E}}}_{n}^{\prime}[g(Y)]$, where $\{Y_{i}^{\prime}\}_{i=1}^{n}$ is an independent copy of $\{Y_{i}\}_{i=1}^{n}$, and thus
$\mathop{\mathbb{E}}_{\{Y_{i}\}_{i=1}^{n}}\sup_{g\in\mathcal{G}}(\mathop{\mathbb{E}}-\widehat{\mathop{\mathbb{E}}}_{n})[g(Y)]$ $\leq\mathop{\mathbb{E}}_{\{Y_{i},Y_{i}^{\prime}\}_{i=1}^{n}}\sup_{g\in\mathcal{G}}(\widehat{\mathop{\mathbb{E}}}_{n}^{\prime}-\widehat{\mathop{\mathbb{E}}}_{n})[g(Y)]$
$=\mathop{\mathbb{E}}_{\{\epsilon_{i}\}_{i=1}^{n}}\mathop{\mathbb{E}}_{\{Y_{i},Y_{i}^{\prime}\}_{i=1}^{n}}\sup_{g\in\mathcal{G}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}(g(Y_{i}^{\prime})-g(Y_{i}))$